LBC & LUCIFER Preparations for Partner Science:
UT 2010 Apr 08

Observer: OKuhn, MPedani (Tucson), DThompson (LBTO)
Telescope Operator: David Gonzalez Huerta
Instrument Support: TShih
Telescope Support: JLittle
SW support: Dan Cox (on call)

Summary

Preparations

Details

22:14 MOS unit subpanel on UI found to be red following mask exchange, but both status windows were green. Initialize cleared it. Mask list not updated yet...asked Tim to do this.

23:16 Mask list updated, IMGUI restarted. Confirmed new names are in UI.

00:04 Configured for N3.75 to check positioning (#0001), looks OK. Also compared to file #0003 from yesterday and they look the same.

01:01 Sky looks completely clear. Tout=2.0, Tglass=-1.5, ambient (next to M1)=4.0 RH=27%

Twilight flats in BrGam and Ks filters
Telescope parked at zenith, facing away from sunset (UT01:52).
  • UTstart 01:07 BrGam+Ks 1x2.0sec Files 0002-0011 DN:12.5k
  • UTstart 01:10 BrGam+clear 1x2.0sec Files 0013-0022 DN:14k
  • UTstart 01:27 BrGam+clear 1x2.0sec Files 0023-0032 DN:11k
  • UTstart 01:41 BrGam+clear 1x2.0sec Files 0033-0042 DN:8k
  • UTstart 01:47 BrGam+clear 1x2.0sec Files 0043-0052 DN:5k
  • UTstart 01:53 BrGam+clear 1x2.0sec Files 0053-0062 DN:2k

last NB=1573 first BB=9000

  • UTstart 01:56 clear+Ks 1x2.0sec Files 0063-0072 DN:9k
  • UTstart 01:59 clear+Ks 1x2.0sec Files 0073-0082 DN:4k

Correct initial collimation/pointing:
  • Initial IE=-48.627, CA=0.381, IA=-384.4
  • Initial collimation looks pretty good, slightly worse than last night but it is not necessary to adjust focus manually
  • At EL=84 corrected IE=-31, CA=7.7
  • At EL=16 corrected IA=-287 (change from last night +97 arcsec!)
  • At EL=82 corrected IE=-36.0, CA=2.1

Collimate on Persson Standard:
02:15 Preset to BS9134 failed...no guide star. Correct pointing nearby: IE=-32.5 CA=12.8
02:21 Preset to BS9134 failed again, no guide star.
02:22 Preset to BS9134 on axis. The star is far enough out of collimation and the seeing may be bad such that the guide star was not found. Allow GCS to collimate here. Seeing is variable 1.5-2.5" and poor ~2" mean, maybe a little better.

02:39 Preset to BS9134 with off-axis guide star

Gather Ks and BrGam illumination correction data (quick sets)
  • script: /home/observer/ObservationScripts/LBTO/box9.DETXY.100as
  • does 10 exposures (repeat at center position at end), +/- 100 arcsec to corners
  • Scripts require you to set up instrument and readout appropriately
  • Data taken around EL=60
  • 02:44 execute box9.DETXY.100as script on BS9134 with off-axis guide star. 10x5.0s int, clear+Ks filters, files 89-98. BS9134 ~ 7000DN/exposure peak above background. 8.5x0.12=1.02 arcsec FWHM
  • 03:04 execute box9.DETXY.100as script, same location. 1x60s normal, BrGam+clear filters, files 100-109. BS9134 ~ 4500DN/exposure peak above background, ~1.3" FWHM.
  • exposure 110 centered, then offset 10.0,10.0 (RADEC REL, so telescope went NE) and...
  • 03:21 execute box9.DETXY.100as script again, 1x60s normal, BrGam+clear filters, files 111-120.
  • offset -20.0,-20.0 (RADEC REL) and...
  • 03:36 execute box9.DETXY.100as script again, 1x60s normal, BrGam+clear filters, files 121-130.
  • Note: offsets are working as expected

Collect field aberration data for Andrew
  • New scripts: /home/LBTO/iraf/ifield_BS9143.cl and /home/LBTO/iraf/ifield_BS9152.cl
  • Correct pointing on nearby pointing star
  • Preset to BS9143 with LUCIFER, allow to collimate
  • Set up in directory /OldRepository/20100408/
  • 04:10 From iraf, execute ifield_BS9143.cl script
    • Script fails to take the agw data. John unreachable. Calling Doug Miller
    • agw_wfsc_image.pro was hard coded for the right side, ask DLM for explanantion
  • 04:47 execute ifield_BS9143.cl script, seems to be working. Completed at 05:13
  • 05:13 execute ifield_BS9143.cl script again. Towards the end the guider images looked out of focus
  • 05:39 execute ifield_BS9143.cl script again, definitely diverging (strong defocus)

06:05 pointto BS9143 gs=18 Tel="GUIDE" PA=30
Run agw_collimate -e 30 -s -f 'LFBG' by hand
Script is giving "Error in irc_sendwavefront"...zernike_send failed. So we were never collimating. What a waste of time!

06:24 sent an active preset with GCS to an off-axis starBS9143 GS 26 PA 30, collimated to less than 200 nm.

06:30 notice that in the M2 PSF GUI there are various yellow values that don't go green, even though the M2 seems to be moving properly and reaching commanded values.

06:33 send guide preset on-axis and take 3 measurements with IDL in "OldRepository/20100408" wfsc 65-67

6:40 Preset to GS 18 PA 110 active preset, converged to ~ 200 nm going on axis to measure WFE.

6:42 sending guide preset on-axis measuring 3 wavefronts, ewfsc 68-70

6:45 preset in active mode to GS 2 PA 230, collimating to < 300 nm

6:52 measuring 3 wavefronts on axis wfsc 71-73

6:55 It seems clear from the three points that we have collimated off axis then measured on axis that there is a sign error in the field aberration correction of Z7 and Z8 now. In other words we are adding the right amount of coma at a given field point instead of subtracting it.

6:56 switching to LBC.

Bias images taken during telescope reconfiguration
  • blind mask+filters
  • Files 0133-0142 10x2.0s int Nexpo=10
  • Files 0143-0147 10x5.0s int Nexpo=10
  • Files 0148-0157 1x60.0s normal Nexpo=10 (telescope slew started at beginning of file 0154, but they all look OK)

Observing Plan (LUCIFER)

  • Collect a pair of 30s LUCIFER images while an in-field offset is taking place
    • Use a bright (pointing) star
    • Check to make sure offsets still happen in a reasonable manner

  • Collect through-focus sequence for Andrew
    • reasonably dense star field.
    • also check the AGw-LUCIFER focus in cold weather for temp term?
    • Offset AGw focus stage by -5.0, -2.5, 0.0, 2.5, 5.0 mm from nominal
    • Allow to collimate, collect 30s total exposure on LUCIFER

  • Check existing pointing model
    • Implement new collimation model if necessary
    • Stable conditions!
    • Take 12 points, with pointing log on...
    • Take full set (56 points) if the first 12 points are off.

  • Collect astrometric field data to check AGw calibration at colder temps?

  • Improve map of AGw patrol field limits at high field angle
    • Requires set up work on a stone field
    • See this plot for extra patrol field area (in blue): AGw_patrolfield_v2.pdf

  • Continue gaining experience with LUCIFER on-sky (e.g. spec acq.)

  • Morning darks

-- DavidThompson - 07 Apr 2010
Topic revision: r8 - 08 Apr 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback